OpenStudy (anonymous):

-6/7x-1/2y=3/7 Match the graph

OpenStudy (anonymous):

Im lost with this one ??! It's my last question

OpenStudy (anonymous):

Is the answer 5/14=3/7???

jimthompson5910 (jim_thompson5910):

It says match, are you given a list of choices?

OpenStudy (anonymous):

Yeah but they are graphs

jimthompson5910 (jim_thompson5910):

are you not able to post them?

jimthompson5910 (jim_thompson5910):

if you can get a hold of the images, you can attach them

OpenStudy (anonymous):

umm i dont think so

OpenStudy (anonymous):

okay i will one sec

jimthompson5910 (jim_thompson5910):

alright

OpenStudy (anonymous):

is that clear enough lol

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

Okayyy. thanks for helping

jimthompson5910 (jim_thompson5910):

yes that works, and it looks great, let me try a few things out for a sec

jimthompson5910 (jim_thompson5910):

The first thing to look for is the slope, how do you find the slope? any ideas?

OpenStudy (anonymous):

Subtract off both sides

OpenStudy (anonymous):

I think its the third one

jimthompson5910 (jim_thompson5910):

solve for y -6/7x-1/2y=3/7 -12x-7y=6 ... multiply every term by the LCD 14 -7y = 12x + 6 y = -12/7x - 6/7 So the slope is -12/7 Which means that it slopes downward, so it cannot be C

OpenStudy (anonymous):

Its B cuz its in the negatives ?? UHHH

OpenStudy (anonymous):

I will medal anyone who gives me the answer :D

jimthompson5910 (jim_thompson5910):

it can't be B either because that slopes upwards

OpenStudy (anonymous):

I've been taking this text for at least 3 hours now :(

jimthompson5910 (jim_thompson5910):

so we've ruled out choices B and C

OpenStudy (anonymous):

Okay so A

jimthompson5910 (jim_thompson5910):

you want something that will slope downwards as x increases from left to right

jimthompson5910 (jim_thompson5910):

does the graph in choice A cut through the y axis below the x-axis?

OpenStudy (anonymous):

Well it can't be D cuz X is a negative

OpenStudy (anonymous):

No A does not

jimthompson5910 (jim_thompson5910):

you nailed it, it's D

OpenStudy (skullpatrol):



OpenStudy (amoodarya):

look at image

jimthompson5910 (jim_thompson5910):

from what I can see, the graph in choice D has a y-intercept below the x-axis A does not, so that rules out choice A

OpenStudy (skullpatrol):